Problem 6 An electron with speed vo 2.37x106 m/s is traveling parallel to and along the same direction as an electric field of magnitudeE7.19x103 N/C. (a) How far will the electron travel before it stops? Submit Answer Tries 0/6 (b) How much time will elapse before it returns to its starting point?
Answer #1

here,

the initial speed of electron , v0 = 2.37 * 10^6 m/s

magnitude of electric feild , E = 7190 N/C

the acceleration of electron , a = charge on electron * E /mass of electron

a = - 1.6 * 10^-19 * 7190 /( 9.1 * 10^-31)

a = - 1.26 * 10^15 m/s^2

a)

let the distance travelled before stopping be s

using third equation of motion

v^2 - v0^2 =2 * a * s

0 - (2.37 * 10^6)^2 = - 2 * 1.26 * 10^15 * s

solving for s

s = 2.23 * 10^-3 m

b)

let the time taken be t

v = v0 + a * t

0 = 2.37 * 10^6 - 1.26 * 10^15 * t

solving for t

t= 1.88 * 10^-9 s

the time taken to return , T = 2 * t

T = 3.76 * 10^-9 s